# Preprint Manuscript

### v3

@@ 2.2.2 Proof Verification (Stage 4)
    Modus Ponens Test (Lemma 4.7): For each triple $(i,j,k)$ with $j,k<i$, the constraint
    $b_{mp,i,j,k}\cdot(f_{i}-f_{j}-f_{k})=0,\quad b_{mp,i,j,k}\in\{0,1\}$ (2.10)
    forces $f_{i}=f_{j}+f_{k}$ when $b_{mp,i,j,k}=1$. The Zeckendorf digits of $f_{i}$ automatically satisfy the non-adjacency constraint by uniqueness (Remark Exposition).
  3. 3.
    Exactly-One Justification (Lemma 4.11): The constraint
    $b_{ax,i}+\sum_{j,k<i}b_{mp,i,j,k}=1$ (2.11)
    ensures that exactly one justification mechanism activates per line.
2.2.3 Degree Management (Stages 5–7)
The proof verification constraints (Stage 4) produce equations of degree $\leq 3$, but merging them into a single equation via sum-of-squares yields degree $6$. We employ two techniques:
  1. 1.
    Guard-Gadgets (Definition 3.6): Each constraint $E(\mathbf{x})=0$ with $\delta(E)\leq 2$ is replaced by the guarded system
    $u\cdot E(\mathbf{x})$ $=0,$ (2.12)
    $u-1-v^{2}$ $=0,$ (2.13)
    where $u,v\in\mathbb{N}$. The second equation forces $u=1+v^{2}\geq 1$, hence $u\neq 0$, so the first equation implies $E(\mathbf{x})=0$. This increases degree to at most $\max(1+2,2)=3$.
  2. 2.
    Jones Factorization (Theorem 5.1): After merging via sum-of-squares (yielding $P_{\text{merged}}$ of degree $6$), we apply recursive monomial shielding. Each monomial $x^{a}y^{b}z^{c}$ with $a+b+c>3$ is factored as $p\cdot q$ where $\delta(p),\delta(q)\leq 3$, introducing auxiliary variables $w$ via constraints $w=p\cdot q$ of degree $\leq 3$. This process (Algorithm 6.3) introduces $O(m\log d)$ variables for a polynomial with $m$ monomials of degree $\leq d$.
2.3 Main Result
Theorem 2.2 (Restatement of Theorem 5.2).
Let $\mathrm{T}$ be a recursively axiomatizable theory with finite axiom set $\mathcal{A}$, and let $G_{T}$ be its Gödel sentence. There exists a polynomial $P(\mathbf{x})\in\mathbb{Z}[\mathbf{x}]$ of degree $\leq 3$ such that:
$T\vdash G_{T}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{m}:P(\mathbf{x})=0.$ (2.14)
The polynomial has $m=O(K^{3}N^{3})$ variables, where $N$ is the proof length and $K=O(\log(\max_{i}f_{i}))$ bounds the Fibonacci representation size.
2.4 On Single Cubic Equations
While our construction produces a system of cubic equations, one might ask whether a single cubic equation could suffice. The following obstruction shows this is impossible.
Proposition 2.3 (Impossibility of Universal Cubic Equation).
There exists no polynomial $P_{\text{univ}}(n,\mathbf{x})\in\mathbb{Z}[n,\mathbf{x}]$ within an $\omega$-consistent formal system of degree $\leq 3$ such that for all Turing machines $M$ (encoded as $n\in\mathbb{N}$):
$M\text{ halts}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{k}:P_{\text{univ}}(n,\mathbf{x})=0.$ (2.15)
Proof sketch.
Assume such $P_{\text{univ}}$ exists. For each machine $M$ encoded as $n$, enumerate solutions $\mathbf{x}\in\mathbb{N}^{k}$ in order of increasing norm $|\mathbf{x}|_{1}=\sum_{i}x_{i}$. By the degree-$3$ constraint, the polynomial $P_{\text{univ}}(n,\mathbf{x})$ has at most $O(k^{3})$ monomials, so checking whether $P_{\text{univ}}(n,\mathbf{x})=0$ takes polynomial time in $|\mathbf{x}|_{1}$.
For halting machines, the smallest solution $\mathbf{x}^{*}$ satisfies $|\mathbf{x}^{*}|_{1}=O(\text{poly}(n))$ (by the polynomial encoding of bounded computations). Thus, we could compute Chaitin [1975] constant.
$\Omega=\sum_{M\text{ halts}}2^{-|M|}$ (2.16)
by enumerating solutions up to polynomial size—contradicting the incomputability of $\Omega$, hence no such $P_{\text{univ}}$ exists. ∎
Remark.
Our construction avoids this obstruction by producing a machine-dependent system $\Phi_{M}(\mathbf{x})$. The proof length $N$ (and hence the variable count $m=O(K^{3}N^{3})$) is not computable from the machine description alone—this non-uniformity is essential. The system encodes "there exists a proof of length NN", not "the shortest proof has length NN".
2.5 Prior Work and Roadmap
We recall:
  1. 1.
    Robinson et al. [1961]: Showed exponential Diophantine relations are listable, reducing Hilbert’s 10th to finding a Diophantine representation of exponentiation. While their construction used higher degrees, our guard-gadget technique demonstrates degree $\leq 3$ suffices for universal representation over $\mathbb{N}$.
  2. 2.
    Jones [1982]: Achieved degree $\leq 4$ via direct MRDP reduction. Our guard-gadget technique (Definition 3.6) reduces this to degree $\leq 3$ by systematically shielding quadratic constraints.
  3. 3.
    Matiyasevich [1970]: Proved undecidability for degree $\leq 4$ (settling the “vanilla” Hilbert’s 10th Problem). Our result shows the degree barrier can be lowered by one, addressing a natural question raised by Jones [1982].
  4. 4.
    Hilbert’s Tenth Problem and its variants for rings, number fields, and related structures remain central questions in mathematics [Alpöge et al., 2025]. However, from a computational perspective, structures like $\mathbb{Z}$ reduce to $\mathbb{N}$ via standard encodings. Mathematical results, see [Poonen, 2009], concern definability complexity in strong foundational theories, which is orthogonal to our focus on bounded provability in weak arithmetics.
The remainder of the paper fills in the technical details:
  1. 1.
    Section 3 (Preliminaries): Establishes encoding tools ($\beta$-function, four-squares, Zeckendorf, guard-gadgets).
  2. 2.
    Section 4 (Construction): Builds the cubic system via proof sequence encoding, axiom testing, and modus ponens verification.
  3. 3.
    Section 5 (Aggregation): Merges the system via sum-of-squares and applies Jones factorization.
  4. 4.
    Section 6 (Computation): Provides algorithmic details for degree reduction.
  5. 5.
    Section 7 (Result): Proves soundness and completeness bidirectionally.
Each stage is constructed to be effective (algorithmically computable), polynomial-time (in the proof parameters), and degree-preserving (maintaining $\delta\leq 3$ after Stage 5).
3 Preliminaries
3.1 The Gödel Beta-Function
The following logic allows encoding arbitrary finite sequences as Diophantine constraints:
Procedure 3.1 (Diophantine $\beta$-Function).
By the work of Gödel [1931] (using the “Chinese remainder”) and as exposited in Matiyasevich [1993], there exists a polynomial $\beta(c,d,i)$ over $\mathbb{N}$ such that for any finite sequence $a_{0},a_{1},\ldots,a_{n}\in\mathbb{N}$, there exist $c,d\in\mathbb{N}$ with
$\beta(c,d,i)=a_{i}\quad\text{for all }i\in\{0,1,\ldots,n\}.$ (3.17)
Explicitly, $\beta(c,d,i)=r_{i}$, where $r_{i}$ is the remainder and $q_{i}\in\mathbb{N}$ is the quotient when dividing $c$ by $(1+(i+1)d)$:
$c=q_{i}\cdot(1+(i+1)d)+r_{i},\quad 0\leq r_{i}<1+(i+1)d.$ (3.18)
Remark.
The “Chinese remainder” allows us to choose $d$ such that the moduli $\{1+(i+1)d:0\leq i\leq n\}$ are pairwise coprime. A suitable choice is
$d=\text{lcm}(1,2,\ldots,n+1),$ (3.19)
which ensures that for any $i\neq j$, the moduli $1+(i+1)d$ and $1+(j+1)d$ differ by (ji)d(j-i)d, and this difference shares no common factor with either modulus, since
$1+(k+1)d\equiv 1\pmod{p}\quad\text{for all primes}\quad p\mid d\quad\text{with}\quad p\leq n+1.$
Hence we construct $c$ via CRT to satisfy $c\equiv a_{i}\pmod{1+(i+1)d}$ for all $i\leq n$. To verify pairwise coprimality with $d=\text{lcm}(1,\ldots,n+1)$: for $i<j\leq n$, we have
$\gcd(1+(i+1)d,1+(j+1)d)\mid\gcd(1+(i+1)d,(j-i)d).$ (3.20)
Since $(j-i)\leq n$ and $d$ is divisible by all integers up to $n+1$, while $1+(i+1)d\equiv 1\pmod{p}$ for any prime $p\mid d$ with $p\leq n+1$, the greatest common divisor equals $1$. This ensures the CRT applies. To express $\beta(c,d,i)=f_{i}$ as polynomial constraints, we introduce variables $r_{i},q_{i}$ and enforce:
$c$ $=q_{i}\cdot(1+(i+1)d)+r_{i},$ (3.21)
$0$ $\leq r_{i}<1+(i+1)d.$ (3.22)
The second challenge, after reducing $\Sigma_{1}$ logic to $\Delta_{0}$ [^4] Foundational reference by Kleene [1943] introducing concepts underlying the arithmetic hierarchy, is encoding inequality polynomially. A naive preliminary approach using $r_{i}=s_{i}^{2}$ to force $r_{i}\geq 0$ failed because: (i) the $\beta$-construction produces remainders $r_{i}=c\bmod(1+(i+1)d)$ that need not be perfect squares (e.g., $r_{i}=2$ is not a square), and (ii) encoding the upper bound $r_{i}<1+(i+1)d$ via $(1+(i+1)d-r_{i}-1)=t_{i}^{2}$ would impose the same impossible requirement. The “four-squares” encoding resolves both issues simultaneously without requiring any specific number to be a perfect square.
Lemma 3.2 (Four-Squares Bound Encoding).
The constraint $0\leq r<m$ over $\mathbb{N}$ is equivalent to the existence of integers $a,b,c,e,\alpha,\beta,\gamma,\delta\in\mathbb{N}$ satisfying:
$r$ $=a^{2}+b^{2}+c^{2}+e^{2},$ (3.23)
$m-1-r$ $=\alpha^{2}+\beta^{2}+\gamma^{2}+\delta^{2}.$ (3.24)
These are polynomial equations of degree $2$.
Proof.
The “Four-Square Theorem” of Lagrange [1770] allows us to say:
Every $n\in\mathbb{N}$ is a sum of four squares.
Constraint (4.48) forces $r\geq 0$. Constraint (4.49) forces $r\leq m-1$, hence $r<m$. Given $r^{*}\in\{0,\ldots,m-1\}$, the four-square decompositions can be computed in polynomial time. By the algorithm of Rabin and Shallit [1986] we decompose an integer $n\in\mathbb{N}$ into four squares
$n=a^{2}+b^{2}+c^{2}+e^{2}$ (3.25)
This ensures the encoding is effective. Computable witnesses exist for any valid bound, and can be constructed in polynomial time (randomized) or quasi-polynomial time (deterministic under ERH). ∎
Remark.
In the completeness proof of Theorem 4.16, the witness construction algorithm (Exposition 4.9) must compute four-square decompositions for each remainder $r_{i}$ and its bound $(i+1)d-1-r_{i}$. By the above algorithm, this can be done in time $O(\log((i+1)d))$ per decomposition, yielding total time $O(N\cdot\log(\max_{i}f_{i}))$ for all $N$ proof lines. This is polynomial in the proof parameters, as required.
3.2 Unique Representation
Remark.
As a “rule of thumb”: A sum satisfies Zeckendorf representation (i.e., uses non-consecutive Fibonacci numbers) when there are “gaps” in the sequence of Fibonacci numbers used. e.g., the tuple $\{F_{6},F_{8},F_{10}\}=\{8,21,55\}$.
Definition 3.3 (Zeckendorf Representation).
Every positive integer $n$ admits a unique representation as a sum of non-consecutive Fibonacci numbers:
$n=\sum_{\kappa\in S}F_{\kappa},$ (3.26)
where $S\subseteq\{2,3,4,\ldots\}$ contains no consecutive integers (i.e., if $\kappa\in S$, then $\kappa+1\notin S$), and $F_{k}$ denotes the $k$-th Fibonacci number with $F_{1}=F_{2}=1$ and $F_{k+2}=F_{k+1}+F_{k}$.
Lemma 3.4 (Polynomial Encoding of Zeckendorf Digits).
For each integer $n$ with Zeckendorf representation involving indices up to $K$, introduce binary variables $d_{2},\ldots,d_{K}\in\{0,1\}$ satisfying:
  1. 1.
    Boolean constraint: $d_{\kappa}-d_{\kappa}^{2}=0$ for all $\kappa\in\{2,\ldots,K\}$,
  2. 2.
    Non-adjacency: $d_{\kappa}\cdot d_{\kappa+1}=0$ for all $\kappa\in\{2,\ldots,K-1\}$,
  3. 3.
    Reconstruction: $n=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{\kappa}$.
These constraints are polynomial of degree at most $2$.
Proof.
The boolean constraint $d_{\kappa}(1-d_{\kappa})=0$ forces $d_{\kappa}\in\{0,1\}$ over $\mathbb{N}$. The product $d_{\kappa}d_{\kappa+1}=0$ enforces non-adjacency. The sum is linear in the $d_{\kappa}$ variables. ∎
Exposition.
The Zeckendorf representation provides an alternative to prime-power encodings for sequences. We use it to enable Fibonacci addition, which can be implemented via digit-wise operations respecting the non-adjacency constraint (see Section 4.3).
3.3 Combined Encoding
Exposition.
For a proof sequence $\langle f_{1},\ldots,f_{N}\rangle$ where each $f_{i}$ is a Gödel number:
  1. 1.
    Encode the sequence using $\beta(c,d,i)=f_{i}$ via constraint (4.47) and Lemma 3.2.
  2. 2.
    Represent each $f_{i}$ in Zeckendorf form using Lemma 3.4.
This dual encoding allows:
  1. 1.
    Variable-length proofs (via existential quantification over $N,c,d$),
  2. 2.
    Arithmetic verification (via Fibonacci addition on Zeckendorf digits).
Regarding the consistency of “Dual Encoding:” Each Gödel number $f_{i}$ is encoded twice:
  1. 1.
    Via the $\beta$-function: $\beta(c,d,i)=f_{i}$ (using variables $c,d,r_{i},q_{i},\ldots$),
  2. 2.
    Via Zeckendorf: $f_{i}=\sum_{\kappa=2}^{K}F_{\kappa}d_{i,\kappa}$ (using digit variables $d_{i,\kappa}$).
These encodings are compatible: for any sequence $\langle f_{1},\ldots,f_{N}\rangle$, witnesses $c^{*},d^{*}$ exist (by Theorem 3.1), and each $f_{i}^{*}$ has a unique Zeckendorf representation (by Definition 3.3). The polynomial system constrains a single variable $f_{i}$ to satisfy both encodings simultaneously, ensuring consistency.
3.4 Summary of Polynomial Constraints
@@
Target check $f_{N}-g_{T}=0$ 1 Linear equality
@@
Target guard $\mathrm{T}\cdot(f_{N}-g_{T})=0$ 2 $\mathrm{T}=1+U^{2}\geq 1$
@@
3.5 Guard-Gadget Framework
Definition 3.5 (Boolean Constraint).
For any variable $b$ intended to represent a boolean value, the constraint
$b-b^{2}=0,\quad b\in\{0,1\}\subseteq\mathbb{N}$ (3.27)
Definition 3.6 (Guarded Constraint).
$u\cdot E(\mathbf{x})$ $=0,$ (3.28)
$u-1-v^{2}$ $=0.$ (3.29)
Proposition 3.7 (Guard Semantics).
The guarded constraint enforces:
  1. 1.
    $u=1+v^{2}\geq 1$ for all solutions (from equation (3.29)),
  2. 2.
    $E(\mathbf{x})=0$ (forced by equation (3.28) since $u\geq 1$).
Proof.
Definition 3.8 (Global Multiplier).
$T=1+U^{2},\quad U\in\mathbb{N}.$ (3.30)
This ensures $T\geq 1$ for all solutions. Optionally, this can be enforced via guarded constraints:
$u_{T}\cdot(T-1-U^{2})$ $=0,$ (3.31)
$u_{T}-1-v_{T}^{2}$ $=0,$ (3.32)
where equation (3.32) forces $u_{T}=1+v_{T}^{2}\geq 1$, and equation (3.31) then enforces $T=1+U^{2}$.
Proposition 3.9 (Selective Activation).
Given a guard $u_{i}$ with $u_{i}=1+v_{i}^{2}$ (from Definition 3.6) and a global multiplier $T\geq 1$, the constraint
$T\cdot(1-u_{i})=0$ (3.33)
forces $u_{i}=1$ (equivalently, $v_{i}=0$).
Proof.
Since $T\geq 1$ and $T\cdot(1-u_{i})=0$, we have $1-u_{i}=0$, hence $u_{i}=1$. From $u_{i}=1+v_{i}^{2}$, this implies $v_{i}=0$. ∎
Exposition.
The selective activation mechanism is used to enforce that specific constraints hold in a solution. For example, in proof verification (Section 4.4), we use it to force exactly one justification (axiom or modus ponens) for each line of a proof. The guard-gadget framework provides a systematic method to enforce boolean constraints without degree inflation (Definition 3.5), ensure all polynomial constraints are non-negative and solvable over $\mathbb{N}$ (Proposition 3.7), and selectively activate constraints via a global multiplier (Proposition 3.9). All gadgets maintain degree $\leq 3$ when applied to linear or quadratic base constraints. This forms the technical foundation for encoding proof verification at cubic degree.
@@ 4 Construction
We now construct a finite system of polynomial equations of degree at most $3$ such that the system has a solution over $\mathbb{N}$ if and only if the Fibonacci variant theory $T_{\text{Fib}}$ proves the target formula $G_{T}$.
@@ 4.1 Proof Sequence Encoding
Exposition.
Constraints (Lemmas 4.14.3) ensure that the sequence $\langle f_{1},\ldots,f_{N}\rangle$ is encoded via $\beta(c,d,i)=f_{i}$ (Theorem 3.1), and each $f_{i}$ has a valid Zeckendorf representation with digits $d_{i,\kappa}$ (Lemma 3.4). The dual encoding is consistent (Exposition). “Why Carry Propagation is Unnecessary.” The key insight is that Zeckendorf representation is unique (Definition 3.3). The constraints enforce $f_{i}=f_{j}+f_{k}$ (arithmetic sum, via Equation (4.59)), and $f_{i}=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa}$ with $d_{i,\kappa}\in\{0,1\}$ and $d_{i,\kappa}\cdot d_{i,\kappa+1}=0$ (via Proposition 4.4). By uniqueness, there exists exactly one choice of $(d_{i,2},\ldots,d_{i,K})$ satisfying the second condition for the value $f_{i}$ determined by the first condition. Hence the constraints force $d_{i,\kappa}$ to be the Zeckendorf digits of $f_{j}+f_{k}$, without explicitly computing the addition “digit-by-digit”. A procedural approach would introduce auxiliary variables $s_{i,j,k,\kappa}$ (sum digits) and $c_{i,j,k,\kappa}$ (carries), computing $s_{\kappa}=d_{j,\kappa}+d_{k,\kappa}+c_{\kappa-1}$ and normalizing to Zeckendorf form. This requires degree-$4$ constraints to handle carry propagation. Our declarative approach avoids this by relying on existential quantification: the solver “guesses” digits $(d_{i,\kappa})$, and uniqueness ensures only the correct guess satisfies all constraints.
@@ 4.2 Combined Proof Encoding
Proof.
Equations (4.47)–(4.49) encode $f_{i}=\beta(c,d,i)$ via the $\beta$-function. Equations (4.50)–(4.52) enforce the Zeckendorf representation of $f_{i}$. Uniqueness of Zeckendorf representation ensures these constraints are consistent. ∎
@@ 4.3 Axiom Membership
Proof.
Immediate from the properties of squares over $\mathbb{N}$. ∎
Exposition.
@@
The key property is that a sum of non-negative terms vanishes if and only if each term vanishes. Over $\mathbb{N}$, this requires $x=a_{j}$ for some $j$. Degree analysis: The sum $\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}$ has degree $2$ (each term is a square of a linear expression). Multiplying by the boolean $b_{ax,i}$ increases the degree to $3$.
@@ 4.4 Modus Ponens Verification
We verify modus ponens by directly enforcing the arithmetic constraint $f_{i}=f_{j}+f_{k}$ for the appropriate triple $(i,j,k)$. The Zeckendorf digit constraints ensure that the digits $d_{i,\kappa}$ represent this sum uniquely, without requiring explicit carry propagation.
@@
The constraint system in Lemma 4.7 correctly encodes modus ponens: Soundness. If $b_{mp,i,j,k}=1$ in a solution, then, equation (4.58) necessitates
$u_{mp,i,j,k}=1+v_{mp,i,j,k}^{2}\geq 1;$ (4.60)
equation (4.59) with $u_{mp,i,j,k}\geq 1$ forces $f_{i}=f_{j}+f_{k}$, while proposition 4.4, the digits $d_{i,\kappa}$ satisfy:
$f_{i}=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa},\quad d_{i,\kappa}\in\{0,1\},\quad d_{i,\kappa}\cdot d_{i,\kappa+1}=0.$ (4.61)
By uniqueness of Zeckendorf representation, the digits $d_{i,\kappa}$ are the unique Zeckendorf digits of $f_{j}+f_{k}$. Completeness. Given a proof where line $i$ is derived from lines $j,k$ via modus ponens (so $f_{i}=f_{j}+f_{k}$ in $\mathbb{N}$): Set $b_{mp,i,j,k}=1$ and $b_{mp,i,j^{\prime},k^{\prime}}=0$ for all $(j^{\prime},k^{\prime})\neq(j,k)$. Compute the Zeckendorf representation of $f_{j}+f_{k}$:
$f_{j}+f_{k}=\sum_{\kappa\in S}F_{\kappa},\quad S\subseteq\{2,\ldots,K\},\;\kappa\in S\implies\kappa+1\notin S.$ (4.62)
We assign $d_{i,\kappa}=\mathbb{1}_{\kappa\in S}$ (indicator function). Set $u_{mp,i,j,k}=1$, $v_{mp,i,j,k}=0$. All constraints (4.57)–(4.59) are satisfied. See [Zeckendorf, 1972] for existence and uniqeness.
Proof.
@@
Encoding the carry logic requires $\delta\geq 4$ to normalize adjacent $1$-s in the binary representation (since Fibonacci addition can produce carries propagating multiple positions). Our key observation: Carry propagation is unnecessary. The constraint $f_{i}=f_{j}+f_{k}$ (linear in $f_{i},f_{j},f_{k}$) combined with the Zeckendorf digit constraints (quadratic in $d_{i,\kappa}$) suffices. The existential quantification $\exists d_{i,\kappa}$ allows the constraint solver to guess the correct digits, which are then verified by:
  1. 1.
@@
The completeness direction of Lemma 4.8 relies on the computability of Zeckendorf representations. Given $n\in\mathbb{N}$, the Zeckendorf digits can be computed by the “greedy algorithm” [^5] An algorithm that makes locally optimal choices at each step, selecting the largest (or best) available option without reconsidering previous decisions. in $O(\log n)$ steps that produces the unique representation satisfying the non-adjacency constraint.:
@@
For verification, given a formal proof $\pi$ of length $N$ with Gödel numbers $f_{1},\ldots,f_{N}$, we can effectively construct a solution $\vec{x}^{*}\in\mathbb{N}^{m}$ by computing $c^{*},d^{*}$ via the $\beta$-function, computing Zeckendorf digits $d_{i,\kappa}^{*}$ for each $f_{i}$ using the greedy algorithm, setting activation variables $b_{ax,i}^{*}$ or $b_{mp,i,j,k}^{*}$ according to the justification of line $i$, and computing all guard variables $u^{*},v^{*}$ to satisfy the positivity constraints. This witness construction is polynomial-time in $N$ and $\max_{i}f_{i}$, demonstrating that the reduction from provability to Diophantine solvability is effective.

### v4

@@ 3.5 Combined encoding
    $E(\mathbf{x})=0$ (forced by equation (3.48) since $u\geq 1$).
Definition 3.11 (Global Multiplier).
$T=1+U^{2},\quad U\in\mathbb{N}.$ (3.50)
This ensures $T\geq 1$ for all solutions. Optionally, this can be enforced via guarded constraints.
@@
Target check $f_{N}-g_{\mathrm{T}}=0$ 1 Linear equality
@@
Target guard $\mathrm{T}\cdot(f_{N}-g_{\mathrm{T}})=0$ 2 $\mathrm{T}=1+U^{2}\geq 1$
@@ 4 Construction
We now construct a finite system of polynomial equations of degree at most $3$ such that the system has a solution over $\mathbb{N}$ if and only if the Fibonacci variant theory $T_{\text{Fib}}$ proves the target formula $G_{\mathrm{T}}$.
@@ 4.2 Combined Proof Encoding
Equations (4.64)–(4.66) encode $f_{i}=\beta(c,d,i)$ via the $\beta$-function. Equations (4.67)–(4.69) enforce the Zeckendorf Representation of $f_{i}$. Uniqueness ensures these constraints are consistent. ∎
@@ 4.3 Axiom Membership
Immediate from the properties of squares over $\mathbb{N}$.
@@ 4.4 Modus Ponens Verification
We verify Modus Ponens by directly enforcing the arithmetic constraint $f_{i}=f_{j}+f_{k}$ for the appropriate triple $(i,j,k)$. The Zeckendorf constraints ensure that the digits $d_{i,\kappa}$ represent this sum uniquely, without requiring explicit carry propagation.
@@
The constraint system in Lemma 4.7 correctly encodes Modus Ponens. Soundness. If
$b_{mp,i,j,k}=1$ (4.77)
in a solution, then (4.75) implies
$u_{mp,i,j,k}=1+v_{mp,i,j,k}^{2}\geq 1$ (4.78)
Hence (4.76) forces
$f_{i}=f_{j}+f_{k}.$ (4.79)
Independently, Proposition 4.4 enforces that $f_{i}$ is reconstructed from Zeckendorf Representation $(d_{i,\kappa})$ with $d_{i,\kappa}\in\{0,1\}$ and $d_{i,\kappa}d_{i,\kappa+1}=0$. By uniqueness, these digits are exactly the Zeckendorf Digits of $f_{j}+f_{k}$.
Completeness. Given a proof in which line $i$ is derived from lines $j,k$ by MP (so $f_{i}=f_{j}+f_{k}$), set $b_{mp,i,j,k}=1$ and all other $b_{mp,i,j^{\prime},k^{\prime}}=0$. Compute the Zeckendorf Digits of $f_{i}$ using Procedure 4.9 and assign $d_{i,\kappa}$ accordingly. Set
$u_{mp,i,j,k}=1$ (4.80)
and
$v_{mp,i,j,k}=0$ (4.81)
Then (4.74)–(4.76) are satisfied.
@@
Encoding the carry logic requires $\delta\geq 4$ to normalize adjacent $1$-s in the binary representation (since Fibonacci addition can produce carries propagating multiple positions). Our key observation: Carry propagation is unnecessary. The constraint $f_{i}=f_{j}+f_{k}$ (linear in $f_{i},f_{j},f_{k}$) combined with the Zeckendorf constraints (quadratic in $d_{i,\kappa}$) suffices. The existential quantification $\exists d_{i,\kappa}$ allows the constraint solver to guess the correct digits, which are then verified by:
  1. (i)
@@
The completeness direction of Lemma 4.8 relies on the computability of “Zeckendorf representation”. Given $n$, the vector is computable by the Greedy Algorithm with an explicit non-adjacency skip. $O(\log n)$ steps that produces the unique representation satisfying the non-adjacency constraint.:
@@
For verification, given a formal proof $\pi$ of length $N$ with Gödel Numbers $f_{1},\ldots,f_{N}$, we can effectively construct a solution
$\vec{x}^{*}\in\mathbb{N}^{m}$ (4.85)
by computing $c^{*},d^{*}$ via the $\beta$-function, computing Zeckendorf Representation $d_{i,\kappa}^{*}$ for each $f_{i}$ using the Greedy Algorithm, setting activation variables $b_{ax,i}^{*}$ or $b_{mp,i,j,k}^{*}$ according to the justification of line $i$, and computing all guard variables $u^{*},v^{*}$ to satisfy the positivity constraints. This witness construction is polynomial-time in $N$ and $\max_{i}f_{i}$, demonstrating that the reduction from provability to Diophantine solvability is effective.